OpenStudy (anonymous):

3 (p - 6 ) = 6 (p + 1) solving proportions?

OpenStudy (anonymous):

6p-18=6p+6 =24

OpenStudy (anonymous):

wait

OpenStudy (jamesj):

No, just use usual algebra. First divide both sides by 3 to get p - 6 = 2(p+1)

OpenStudy (anonymous):

i screwed up

OpenStudy (jamesj):

That means multiplying out the 2(p+1) p - 6 = 2p + 2 Hence - 6 = p + 2 I.e., -8 = p

OpenStudy (anonymous):

3p-18=6p+6 -3p=24 p=-8

OpenStudy (anonymous):

what i don'tundersand

OpenStudy (anonymous):

Zluboldt i understand you but not JamesJ I don't understand yours

OpenStudy (anonymous):

yeah, i don't know what he was doing

OpenStudy (jamesj):

3 (p - 6 ) = 6 (p + 1) First divide both sides by 3 to get p - 6 = 2(p+1) Is that clear?

OpenStudy (anonymous):

oh, i just did that at the last step

OpenStudy (anonymous):

no you are not spouse to divide

OpenStudy (anonymous):

i like my way better

OpenStudy (anonymous):

me to

OpenStudy (jamesj):

Fine. Both methods are correct.